OpenStudy (anonymous):

Translate each sentence into an equation. 1. Three more than four times a number is equal to twelve. 2. The sum of a number and two is twice the number. 3. The quotient of a number and twelve is the sum of eight and the number. 4. Five more than three times a number is equal to eight. 5. The sum of a number and six is seven times the number. 6. The quotient of a number and six is equal to nine. 7. Seven less than a number is two. 8. Twenty-seven is the product of three and a number. 9. Twelve less than six times a number is equal to three times the number.

OpenStudy (phi):

It would be *very* useful if you could answer these on your own. A few minutes of work on these problems is worth a whole lot later on...

OpenStudy (anonymous):

1. 4x+3=12 2. x+2 = 2x 3. x/12 =y+8 4. 3x+5= 8 5. x+6 =7x 6. x/6 =9 7. x-7 = 2 8. 3 * x = 27 9. 6x-12 = 3x

OpenStudy (anonymous):

are the answers correct ?

OpenStudy (anonymous):

ya wow u are awesome! ;) what grade are you in?

OpenStudy (anonymous):

i'm am in class 12

OpenStudy (anonymous):

and u??

OpenStudy (anonymous):

in 9th :)

OpenStudy (anonymous):

which country?????

OpenStudy (anonymous):

USA u?

OpenStudy (anonymous):

INDIA